Use scientific notation to multiply the following two numbers. Express the answer using scientific notation; retain all nonzero decimal places.
(2.6 · 10^2) · (7.3 · 10^3)

Found 2 solutions by mananth, Darksinz29:
Answer by mananth(16949)

(2.6 · 10^2) · (7.3 · 10^3)
= (2.6*7.2*10^2*10^3)
=(18.72*10^5)
=1.872*10^6
...
m.ananth@hotmail.ca

Answer by Darksinz29(2)

2.6*7.3*10^2*10^3=
18.98 * 10^5
1.898 * 10^6